OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 6 + 14 + . . . + 2n = 2(n+1) - 2

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

r u sure?

OpenStudy (helder_edwin):

if n=1 then 2n=2 (the first term) but if n=2 then 2n=2*2=4 not the second term !!!

OpenStudy (anonymous):

So what would we say for that?

OpenStudy (helder_edwin):

it doesn't ask u to prove it. does it??

OpenStudy (anonymous):

Well it just says to write the statement.

OpenStudy (helder_edwin):

OK. then let's ignore that issue then. but don't forget to do what i did whenever u r asked to prove this kind of stuff. OK?

OpenStudy (anonymous):

Okay I won't

OpenStudy (helder_edwin):

u have \[ \large P(n):\quad 2+6+14+\dots+2n=2(n+1)-2 \]

OpenStudy (helder_edwin):

do u know what P(1) means ??

OpenStudy (anonymous):

Yes I do

OpenStudy (helder_edwin):

so do it. write what u think the answer is

OpenStudy (anonymous):

so the right side would be 2?

OpenStudy (helder_edwin):

yes. and the left side?

OpenStudy (anonymous):

2(1)=2?

OpenStudy (helder_edwin):

yes.

OpenStudy (helder_edwin):

so \[ \large P(1): 2=2 \]

OpenStudy (helder_edwin):

what about P(k) ??

OpenStudy (anonymous):

What would that be?

OpenStudy (helder_edwin):

it means, replace EVERY n by k

OpenStudy (anonymous):

So it would be P(k)= P(1)?

OpenStudy (helder_edwin):

no. in the original problem, just replace every \(n\) with a \(k\).

OpenStudy (anonymous):

oh so then after that, we do Pk+1= P(1+1)?

OpenStudy (helder_edwin):

no. it is quite the same P(k+1) means "replace every \(n\) with \(k+1\). in the original problem.

OpenStudy (helder_edwin):

so. for instance \[ \large P(k):2+6+14+\dots+2k=2(k+1)-2 \]

OpenStudy (helder_edwin):

try writting P(k+1)

OpenStudy (anonymous):

2(k+1)? I'm not quite sure

OpenStudy (helder_edwin):

no. the whole thing. look at what i wrote for P(k) !!

OpenStudy (anonymous):

P(k+1)=2 + 6 + 14 + . . . + 2k = 2(k+1) - 2?

OpenStudy (helder_edwin):

no. every n has to be replaced by k+1. so for instance, if u had 7n+3 it would then become 7(k+1)+3 got it?

OpenStudy (anonymous):

ohh yeah okay. So what would I basically write for the answer though?

OpenStudy (helder_edwin):

well \[ \large P(k+1):2+6+14+\dots+2(k+1)=2[(k+1)+1]-2 \]

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

And that would be it?

OpenStudy (helder_edwin):

yes.

OpenStudy (anonymous):

Great :) thank you so much again! I really appreciate it

OpenStudy (helder_edwin):

u r very welcome.